# Attributed-graphs kernel implementation using local detuning of neutral-atoms Rydberg Hamiltonian

### v1

@@ IV Results
Figure 4: Relative rank of the Gram matrix $\mathrm{rank}(K)/N_{\mathrm{data}}$ associated with the GDQC kernel $\kappa_{\mathrm{GDQC}}$ (13) versus the F1 score of the corresponding kernel-based SVM for varying binning number $N_{\mathrm{bins}}^{C}$. The quantities are plotted as a function of the time steps across the entire time evolution with $t_{\mathrm{max}}=1\,\mathrm{\mu s}$. (a)-(d) corresponds to the dataset MUTAG; (e)-(h) corresponds to PTC_FM. The identifier $\mathrm{(loc)}$ $\mathrm{(glob)}$] imply that the involved time evolution is governed by the Hamiltonian $\hat{\mathcal{H}}_{\mathrm{loc}}^{\mathcal{G}}$ (10) $\hat{\mathcal{H}}_{\mathrm{glob}}^{\mathcal{G}}$ (9)], whereby $\hat{\mathcal{H}}_{\mathrm{loc}}^{\mathcal{G}}$ has node features embedded. For comparison, the black dashed line with x-markers represents the relative rank associated with the QEK kernel $\kappa_{\mathrm{QEK}}$ (12), which does not depend on binning.
@@
maximum evolution time: $t_{\mathrm{max}}=1\,\mathrm{\mu s}$
$\mathrm{QEK}^{\mathrm{(loc)}}_{\bm{\times}}$ {0.2, 0.75, 0.95} 86.75 $\pm$ 5.67 {0.1, 0.4, 0.9} 63.33 $\pm$ 5.90
{0.15, 0.25, 0.4, 0.65, 0.95} 87.41 $\pm$ 6.2 {0.05, 0.2, 0.4, 0.5, 0.95} 63.89 $\pm$ 6.15
$\mathrm{QEK}^{\mathrm{(loc)}}_{\bm{+}}$ {0.2, 0.75, 0.9} 87.08 $\pm$ 5.81 {0.2, 0.7, 0.95} 62.40 $\pm$ 5.77
{0.05, 0.1, 0.25, 0.75, 0.9} 86.80 $\pm$ 5.53 {0.25, 0.4, 0.7, 0.95, 1.} 62.35 $\pm$ 5.70
$\mathrm{QEK}^{\mathrm{(glob)}}_{\bm{\times}}$ {0.05 0.15 0.2} 85.7 $\pm$ 5.44 {0.05 0.35 0.4} 62.42 $\pm$ 5.66
{0.1 0.25 0.8 0.85 0.9} 85.01 $\pm$ 6.2 {0.05 0.1 0.2 0.35 0.4} 60.77 $\pm$ 6.02
$\mathrm{QEK}^{\mathrm{(glob)}}_{\bm{+}}$ {0.1 0.2 0.8} 84.98 $\pm$ 5.71 {0.2 0.35 0.4} 62.74 $\pm$ 5.12
{0.05 0.1 0.2 0.8 0.85} 84.96 $\pm$ 5.47 {0.05 0.2 0.25 0.35 0.4} 61.48 $\pm$ 5.13
$\mathrm{GDQC}^{\mathrm{(loc)}}_{\bm{\times}}$ {0.05 0.15 0.8} 85.92 $\pm$ 5.95 {0.45 0.5 0.9} 62.98 $\pm$ 5.31
{0.05 0.1 0.15 0.25 0.35} 85.29 $\pm$ 5.57 {0.3 0.45 0.5 0.55 0.9} 63.71 $\pm$ 5.26
$\mathrm{GDQC}^{\mathrm{(loc)}}_{\bm{+}}$ {0.1 0.35 0.8} 86.49 $\pm$ 5.22 {0.25 0.35 0.45} 61.96 $\pm$ 5.87
{0.1 0.35 0.8 0.85 0.9} 86.0 $\pm$ 5.67 {0.1 0.25 0.35 0.55 0.85} 61.83 $\pm$ 5.55
$\mathrm{GDQC}^{\mathrm{(glob)}}_{\bm{\times}}$ {0.1 0.35 0.8} 87.77 $\pm$ 4.72 {0.05 0.25 0.45} 59.73 $\pm$ 5.54
{0.05 0.15 0.2 0.35 0.8} 87.06 $\pm$ 5.18 {0.15 0.25 0.45 0.6 0.9} 59.07 $\pm$ 5.32
$\mathrm{GDQC}^{\mathrm{(glob)}}_{\bm{+}}$ {0.35 0.65 0.8} 88.06 $\pm$ 5.0 {0.7 0.95 1.0} 57.96 $\pm$ 4.4
{0.05 0.25 0.55 0.8 0.95} 88.0 $\pm$ 5.1 {0.15 0.6 0.7 0.95 1.0} 58.08 $\pm$ 5.1
maximum evolution time: $t_{\mathrm{max}}=2\,\mathrm{\mu s}$
$\mathrm{QEK}^{\mathrm{(loc)}}_{\bm{\times}}$ {0.2 1.2 1.6} 86.83 $\pm$ 5.6 {0.1 0.4 1.0} 63.12 $\pm$ 6.0
{0.2 0.3 0.5 1.2 1.6} 86.26 $\pm$ 5.6 {0.2 0.4 0.7 1.0 1.1} 63.17 $\pm$ 5.5
$\mathrm{QEK}^{\mathrm{(loc)}}_{\bm{+}}$ {0.2 0.6 0.9} 85.8 $\pm$ 5.9 {0.3 1.3 1.8} 62.29 $\pm$ 5.9
{0.2 0.3 0.7 1.2 1.6} 85.67 $\pm$ 5.8 {0.3 1.3 1.7 1.8 1.9} 62.26 $\pm$ 5.9
$\mathrm{QEK}^{\mathrm{(glob)}}_{\bm{\times}}$ {0.1 0.2 1.5} 86.53 $\pm$ 4.8 {0.3 1.5 1.6} 61.76 $\pm$ 4.9
{0.1, 0.2, 1.3, 1.5, 2.0} 86.05 $\pm$ 5.16 {0.4 1.4 1.5 1.6 1.8} 61.94 $\pm$ 5.5
$\mathrm{QEK}^{\mathrm{(glob)}}_{\bm{+}}$ {0.2 1.5 2.0} 86.33 $\pm$ 5.1 {0.3 1.5 1.6} 61.2 $\pm$ 5.1
{0.2 0.8 1.3 1.5 1.7} 85.68 $\pm$ 5.5 {0.2 0.3 1.3 1.4 1.5} 61.99 $\pm$ 5.0
$\mathrm{GDQC}^{\mathrm{(loc)}}_{\bm{\times}}$ {0.1 0.8 1.4} 86.82 $\pm$ 5.6 {0.5 1.3 1.6} 62.87 $\pm$ 6.0
{0.5 0.8 1.1 1.2 1.5} 86.88 $\pm$ 4.9 {0.4 0.5 0.9 1.1 1.6} 65.38 $\pm$ 5.4
$\mathrm{GDQC}^{\mathrm{(loc)}}_{\bm{+}}$ {0.6 0.8 1.4} 86.78 $\pm$ 5.2 {1.3 1.6 1.7} 61.04 $\pm$ 5.7
{0.2 0.8 1.1 1.2 1.4} 87.3 $\pm$ 5.1 {0.5 0.8 1.3 1.6 1.7} 62.18 $\pm$ 5.2
$\mathrm{GDQC}^{\mathrm{(glob)}}_{\bm{\times}}$ {0.3 0.8 1.3} 87.11 $\pm$ 4.9 {1.0 1.3 1.6} 60.33 $\pm$ 5.5
{0.1 0.7 0.8 1.2 1.3} 87.32 $\pm$ 5.0 {0.1 0.2 0.6 1.0 1.6} 58.86 $\pm$ 5.2
$\mathrm{GDQC}^{\mathrm{(glob)}}_{\bm{+}}$ {0.5 0.7 0.8} 86.51 $\pm$ 5.2 {0.8 1.6 1.8} 59.12 $\pm$ 5.0
{0.3 0.5 0.8 1.3 1.6} 87.18 $\pm$ 5.1 {0.2 0.8 1.3 1.6 1.8} 59.22 $\pm$ 5.8
IV.1 Binning-size dependence of GDQC
@@
In this part we study the impact of the chosen binning size $\varDelta_{C}=1/N_{\mathrm{bins}}^{C}$ which is applied to the correlation matrix $C^{\mathcal{G}}_{i,j}\left(t\right)$ within the GDQC framework, cf. Subsec. III.2.
For this purpose, we vary the number of bins $N_{\mathrm{bins}}^{C}$ and we study two quantities providing complementary information. First, we compute the F1 score showcasing the actual performance of the respective GDQC model. Second, we compare it with the rank of the Gram matrix $K$ where $K_{\mu,\mu^{\prime}}=\kappa_{\mathrm{GDQC}}\left(\mathcal{G}_{\mu},\mathcal{G}_{\mu^{\prime}}\right)$ with $\mu,\mu^{\prime}\in\left\{1,2,\dots,N_{\mathrm{data}}\right\}$ and the size of the dataset $N_{\mathrm{data}}$. The rank of a matrix indicates how many of its constituent vectors are linearly independent. As such the rank of the Gram matrix $\mathrm{rank}\left(K\right)$ provides insights into the expressiveness of the kernel; the more linearly-independent GDQC feature vectors $\bm{\chi}\left[C^{\mathcal{G}_{\mu}}\left(t\right),D^{\mathcal{G}_{\mu}}\right]$, the more expressive the kernel. The two quantities, the relative rank $\mathrm{rank}\left(K\right)/N_{\mathrm{data}}$ and the F1 score, are plotted in Fig. 4 for varying number of bins $N_{\mathrm{bins}}^{C}$ and across the entire time evolution with $t_{\mathrm{max}}=1\,\mathrm{\mu s}$.
The first thing to note is that increasing the number of bins $N_{\mathrm{bins}}^{C}$ does improve the expressiveness of the GDQC kernel $\kappa_{\mathrm{GDQC}}$ (13) as the rank of the Gram matrix increases significantly. Beyond a certain number of bins, the relative rank saturates. Interestingly, the saturation for the local-detuning case $\mathrm{GDQC}^{\mathrm{(loc)}}$ reaches almost $1$ [panels (c),(g)] whereas the global-detuning case $\mathrm{GDQC}^{\mathrm{(glob)}}$ saturates at a smaller value [panels (a),(e)]. This difference can be traced back to the qualitative difference between both kernel implementations: the local-detuning kernel being more expressive. In this particular case, there are molecules in the datasets which share the same atomic placements (the non-attributed underlying graph) but whose atomic constituents (the node features) are different. The local-detuning kernel $\kappa_{\mathrm{GDQC}}^{\mathrm{loc}}$ is able to differentiate such molecules whereas the global-detuning kernel $\kappa_{\mathrm{GDQC}}^{\mathrm{loc}}$ is not, resulting in a lower Gram-matrix rank for the latter. The tiny difference between $1$ and the local-detuning plateaus in panels (c) and (g) can be traced back to a few molecules which only differ by the presence of hydrogen atoms–which are omitted in the embedded graphs. The second main observation is that increasing the expressiveness of the kernel does not increase the performances of the SVM model. Indeed in both data sets, the F1 score is comparable when computing the $\kappa_{\mathrm{GDQC}}$ with $N_{\mathrm{bins}}^{C}=10$ and with $N_{\mathrm{bins}}^{C}=1000$, even when the rank of the Gram matrix is 4 times larger.
At last we compare the expressiveness of the two quantum-feature kernels studied in this work, the GDQC kernel $\kappa_{\mathrm{GDQC}}$ (13) and the QEK kernel $\kappa_{\mathrm{QEK}}$ (12). To this end, we added the rank of the Gram matrix associated with QEK to Figs. 4(a),(c),(e),(g), shown as the black dashed line with x markers. Recall that there is no binning dependence in the QEK algorithm. Interestingly, the rank of $\kappa_{\mathrm{QEK}}$ is as low as it is for $\kappa_{\mathrm{GDQC}}$ with only $N_{\mathrm{bins}}^{C}=10$. One conjecture is that the proposed GDQC kernel can be chosen such that its expressiveness is significantly larger compared than what it is for the QEK kernel. Nonetheless, regardless the expressiveness the corresponding F1 scores are very similar, cf. Fig. 3. This finding supports the hypothesis that the performance of a kernel-based machine learning model is not only influenced by the expressiveness of the involved kernel, but also by its capacity to associate the relevant "similarity scores" to the corresponding data points.

### v2

@@ IV Results
Figure 4: Relative rank of the Gram matrix $\mathrm{rank}(K)/N_{\mathrm{data}}$ (23) associated with the $\mathrm{GDQC}$ kernel $\kappa_{\mathrm{GDQC}}$ (17) versus the F1 score of the corresponding kernel-based SVM for varying binning number $N_{\mathrm{bins}}^{C}$. The quantities are plotted as a function of the time steps across the entire time evolution with $t_{\mathrm{max}}=1\,\mathrm{\mu s}$. (a)-(d) corresponds to the dataset MUTAG; (e)-(h) corresponds to PTC_FM. The identifier $\mathrm{(loc)}$ $\mathrm{(glob)}$] imply that the involved time evolution is governed by the Hamiltonian $\hat{\mathcal{H}}_{\mathrm{loc}}^{\mathcal{G}}$ (10) $\hat{\mathcal{H}}_{\mathrm{glob}}^{\mathcal{G}}$ (9)], whereby $\hat{\mathcal{H}}_{\mathrm{loc}}^{\mathcal{G}}$ has node features embedded. For comparison, the black dashed curve with x-markers represents the data associated with the $\mathrm{QEK}$ kernel $\kappa_{\mathrm{QEK}}$ (15), which does not depend on binning. The horizontal gray dotted line corresponds to the theoretical relative rank value computed from the number of distinct graph-isomorphism classes.
@@
maximum evolution time: $t_{\mathrm{max}}=1\,\mathrm{\mu s}$ $\mathrm{QEK}^{\mathrm{(loc)}}$ 1 84.52 $\pm$ 5.34 0.5 64.11 $\pm$ 4.75
$\mathrm{QEK}^{\mathrm{(loc)}}_{\bm{\times}}$ {0.24, 0.44, 0.74} 87.05 $\pm$ 6.04 {0.14, 0.4, 1.} 66.54 $\pm$ 5.02
{0.24, 0.3, 0.34, 0.74, 0.9} 87.61 $\pm$ 5.55 {0.1, 0.84, 0.9, 0.94, 1.} 66.66 $\pm$ 4.53
$\mathrm{QEK}^{\mathrm{(loc)}}_{\bm{+}}$ {0.24, 0.44, 0.64} 86.61 $\pm$ 5.97 {0.4, 0.94, 1} 67.34 $\pm$ 4.05
{0.14, 0.24, 0.34, 0.44, 0.64} 86.89 $\pm$ 5.56 {0.04, 0.14, 0.34, 0.4, 1.} 67.29 $\pm$ 4.17
$\mathrm{QEK}^{\mathrm{(glob)}}$ 0.1 84.62 $\pm$ 5.48 0.44 60.62 $\pm$ 5.16
$\mathrm{QEK}^{\mathrm{(glob)}}_{\bm{\times}}$ {0.14, 0.64, 0.7} 87.04 $\pm$ 4.46 {0.2, 0.24, 1.} 61.27 $\pm$ 5.05
{0.1, 0.24, 0.64, 0.7, 0.74} 87.35 $\pm$ 5.15 {0.14, 0.2, 0.34, 0.4, 1.} 62.6 $\pm$ 5.79
$\mathrm{QEK}^{\mathrm{(glob)}}_{\bm{+}}$ {0.2, 0.64, 0.7} 87.47 $\pm$ 4.9 {0.24, 0.34, 1} 60.78 $\pm$ 4.72
{0.1, 0.24, 0.64, 0.7, 0.74} 88.21 $\pm$ 5.92 {0.2, 0.24, 0.34, 0.4, 1.} 61.19 $\pm$ 4.89
$\mathrm{GDQC}^{\mathrm{(loc)}}$ 0.5 89.29 $\pm$ 4.89 1 60.87 $\pm$ 5.49
$\mathrm{GDQC}^{\mathrm{(loc)}}_{\bm{\times}}$ {0.5, 0.74, 0.94} 89.58 $\pm$ 4.81 {0.4, 0.84, 1.} 66.41 $\pm$ 5.70
{0.24, 0.5, 0.64, 0.7, 0.94} 90.09 $\pm$ 5.05 {0.24, 0.34, 0.4, 0.84, 1.} 66.90 $\pm$ 4.58
$\mathrm{GDQC}^{\mathrm{(loc)}}_{\bm{+}}$ {0.04, 0.5, 0.64} 88.74 $\pm$ 5.69 {0.34, 0.84, 1.} 64.09 $\pm$ 4.97
{0.3, 0.54, 0.6, 0.74, 0.9} 88.81 $\pm$ 5.55 {0.3, 0.5, 0.64, 0.84, 1.} 65.69 $\pm$ 4.86
$\mathrm{GDQC}^{\mathrm{(glob)}}$ 0.54 88.53 $\pm$ 4.33 0.3 59.61 $\pm$ 5.57
$\mathrm{GDQC}^{\mathrm{(glob)}}_{\bm{\times}}$ {0.24, 0.4, 0.74} 88.95 $\pm$ 5.50 {0.04, 0.1, 0.24} 61.19 $\pm$ 5.44
{0.04, 0.1, 0.24, 0.4, 0.54} 88.82 $\pm$ 5.07 {0.2, 0.6, 0.64, 0.8, 0.84} 61.25 $\pm$ 5.61
$\mathrm{GDQC}^{\mathrm{(glob)}}_{\bm{+}}$ {0.24, 0.4, 0.54} 88.35 $\pm$ 5.18 {0.24, 0.3, 0.84} 59.32 $\pm$ 4.84
{0.04, 0.2, 0.24, 0.4, 0.54} 88.60 $\pm$ 5.27 {0.24, 0.4, 0.6, 0.64, 0.8} 59.43 $\pm$ 4.39
maximum evolution time: $t_{\mathrm{max}}=2\,\mathrm{\mu s}$ $\mathrm{QEK}^{\mathrm{(loc)}}$ 2 86.19 $\pm$ 5.03 0.5 64.11 $\pm$ 4.75
$\mathrm{QEK}^{\mathrm{(loc)}}_{\bm{\times}}$ {1.3, 1.4, 1.8} 88.92 $\pm$ 4.59 {0.1, 1.1, 1.6} 68.32 $\pm$ 5.44
{0.1, 1.1, 1.3, 1.4, 2} 88.78 $\pm$ 4.89 {0.4, 1.1, 1.3, 1.5, 1.6} 69.28 $\pm$ 5.28
$\mathrm{QEK}^{\mathrm{(loc)}}_{\bm{+}}$ {1.4, 1.6, 1.7} 88.16 $\pm$ 5.99 {0.2, 1.1, 1.5} 68.65 $\pm$ 5.96
{0.1, 0.7, 1.4, 1.5, 1.8} 89.26 $\pm$ 5.40 {1., 1.1, 1.3, 1.5, 1.9} 69.55 $\pm$ 4.98
$\mathrm{QEK}^{\mathrm{(glob)}}$ 2 85.8 $\pm$ 5.81 0.44 60.62 $\pm$ 5.16
$\mathrm{QEK}^{\mathrm{(glob)}}_{\bm{\times}}$ {0.3, 0.5, 1.5} 88.57 $\pm$ 4.03 {1., 1.3, 1.5} 62.56 $\pm$ 5.87
{0.2, 0.5, 1.1, 1.4, 1.5} 88.96 $\pm$ 4.98 {0.3, 1.1, 1.2, 1.3, 1.5} 63.35 $\pm$ 4.78
$\mathrm{QEK}^{\mathrm{(glob)}}_{\bm{+}}$ {0.2, 1.5, 2.} 87.63 $\pm$ 5.58 {0.2, 0.4, 1.9} 62.66 $\pm$ 4.92
{0.2, 0.5, 0.7, 1.4, 1.7} 89.45 $\pm$ 4.02 {0.3, 0.4, 1., 1.8, 1.9} 62.22 $\pm$ 5.54
$\mathrm{GDQC}^{\mathrm{(loc)}}$ 0.5 89.29 $\pm$ 4.89 1 60.87 $\pm$ 5.49
$\mathrm{GDQC}^{\mathrm{(loc)}}_{\bm{\times}}$ {1.3, 1.4, 1.7} 90.25 $\pm$ 5.02 {0.04, 0.84, 1.} 66.41 $\pm$ 5.70
{0.1, 0.5, 1.3, 1.4, 1.7} 90.83 $\pm$ 5.56 {0.4, 1.1, 1.5, 1.6, 1.7} 66.51 $\pm$ 4.82
$\mathrm{GDQC}^{\mathrm{(loc)}}_{\bm{+}}$ {0.5, 1.3, 1.9} 89.28 $\pm$ 4.21 {0.3, 1., 1.9} 65.48 $\pm$ 4.21
{0.5, 0.6, 1.1, 1.3, 1.8} 89.35 $\pm$ 4.97 {0.1, 0.3, 1., 1.3, 1.6} 65.86 $\pm$ 5.26
$\mathrm{GDQC}^{\mathrm{(glob)}}$ 0.8 88.53 $\pm$ 4.33 0.3 59.61 $\pm$ 5.57
$\mathrm{GDQC}^{\mathrm{(glob)}}_{\bm{\times}}$ {0.8, 1.2, 1.3} 88.65 $\pm$ 5.56 {0.6, 1.2, 1.7} 60.4 $\pm$ 4.85
{0.5, 1.3, 1.4, 1.7, 1.9} 89.10 $\pm$ 5.55 {0.3, 0.6, 1.1, 1.4, 1.7} 61.18 $\pm$ 5.66
$\mathrm{GDQC}^{\mathrm{(glob)}}_{\bm{+}}$ {0.8, 1.2, 1.3} 88.84 $\pm$ 5.42 {0.1, 1.4, 1.5} 59.41 $\pm$ 5.23
{0.8, 0.9, 1.2, 1.3, 1.9} 88.84 $\pm$ 5.48 {0.3, 0.6, 1., 1.1, 1.7} 59.83 $\pm$ 6.02
IV.1 Binning-size dependence, graph distinguishability and limitations
@@
In this part, we study the impact of the correlation-binning resolution $\varDelta_{C}=1/N_{\mathrm{bins}}^{C}$ applied to the correlation matrix $C^{\mathcal{G}}_{i,j}(t)$ within the $\mathrm{GDQC}$ framework (cf. Subsec. III.3).
To this end, we vary the number of correlation bins $N_{\mathrm{bins}}^{C}$ and track two complementary quantities. First, we report the weighted F1 score, which captures the predictive performance of the corresponding $\mathrm{GDQC}$ model. Second, we compute the rank of the associated Gram matrix $K\in\mathbb{R}^{N_{\mathrm{data}}\times N_{\mathrm{data}}}$ with entries
$K_{\mu,\mu^{\prime}}$ $=\kappa_{\mathrm{GDQC}}\!\left(\mathcal{G}_{\mu},\mathcal{G}_{\mu^{\prime}}\right),$ $\mu,\mu^{\prime}$ $\in\{1,\dots,N_{\mathrm{data}}\},$ (23)
where $N_{\mathrm{data}}$ denotes the dataset size, and the $\mathrm{GDQC}$ kernel $\kappa_{\mathrm{GDQC}}$ is defined in Eq. (17). We estimate $\mathrm{rank}(K)$ by counting the number of singular values above a numerical threshold [37, 34]. The normalized quantity $\mathrm{rank}(K)/N_{\mathrm{data}}\in[0,1]$ serves as a proxy for the effective dimensionality of the kernel representation on the sampled dataset, and hence for the degree to which the kernel induces non-degenerate (linearly independent) feature representations. Figure 4 reports both the relative rank $\mathrm{rank}\left(K\right)/N_{\mathrm{data}}$ and the weighted F1 score for varying $N_{\mathrm{bins}}^{C}$, across the time evolution up to $t_{\mathrm{max}}=1\,\mathrm{\mu s}$.
We observe that increasing the number of correlation bins $N_{\mathrm{bins}}^{C}$ increases the rank of the $\mathrm{GDQC}$ Gram matrix, i.e., the effective dimensionality of $\kappa_{\mathrm{GDQC}}$ in Eq. (17). This suggests that, on the present datasets, finer binning yields a less degenerate similarity structure and allows the kernel to separate a larger number of nearby graphs. Beyond a certain $N_{\mathrm{bins}}^{C}$, the relative rank $\mathrm{rank}\left(K\right)/N_{\mathrm{data}}$ saturates. The saturation level is close to $1$ for the local-detuning variant $\mathrm{GDQC}^{(\mathrm{loc})}$ (panels (c),(g)), whereas it remains substantially smaller for the global-detuning variant $\mathrm{GDQC}^{(\mathrm{glob})}$ (panels (a),(e)), consistent with local detuning inducing a richer empirical similarity structure than global detuning on these datasets.
To interpret these rank values, we also compute an idealized reference rank from the number and cardinalities of graph isomorphism classes (shown as the gray dotted line in Fig. 4), where isomorphism classes are determined using the VF2 algorithm [13]. This reference depends on whether node attributes are included in the isomorphism test: ignoring attributes yields fewer distinct classes and hence a smaller reference rank than the attribute-aware case, which is expected since some molecules differ only by atom and bond types. Notably, for a large number of correlation bins (fine binning), the empirical rank of the $\mathrm{GDQC}$ Gram matrix can exceed this combinatorial reference value. This reflects a limitation of our current pipeline. Indeed, the embedding computation involves stochasticity through the random initialization of atomic positions. As a result, repeated embeddings of the same graph do not necessarily converge to identical positions, which can lead to slightly different quantum features. This effect is present for both $\mathrm{GDQC}$ and $\mathrm{QEK}$, but $\mathrm{GDQC}$ is more sensitive due to the binning step: under fine binning, small variations in correlation values can be amplified. For moderate binning, i.e., $N_{\mathrm{bins}}^{C}\lesssim 100$, the induced RKHS representations of two isomorphic graphs remain very close. For normalized kernels, this can be quantified via
$\left\|\bm{\chi}(\mathcal{G}_{\mu})-\bm{\chi}(\mathcal{G}_{\mu^{\prime}})\right\|^{2}$ $=2-2K_{\mu,\mu^{\prime}}\ll 1.$ (24)
This suggests that the perturbation induced by stochastic embeddings is limited in that regime. Quantifying its impact on downstream SVM performance would require an explicit robustness analysis across random initializations, which we leave for future work.
For completeness, we perform the same Gram-rank analysis for the $\mathrm{QEK}$ kernel $\kappa_{\mathrm{QEK}}$ (15) and include the corresponding results in Fig. 4 (black dashed line with $\times$ markers). Recall that $\mathrm{QEK}$ has no binning dependence. Interestingly, the rank of $\kappa_{\mathrm{QEK}}$ is comparable to that of $\kappa_{\mathrm{GDQC}}$ at only $N_{\mathrm{bins}}^{C}=10$. In the present experimental setting, $\mathrm{GDQC}$ therefore offers a broader tunable range of empirical graph distinguishability through the binning parameter. However, this increased distinguishability does not yield a clear improvement in weighted F1 score in our experiments (cf. Fig. 3).
This finding highlights that graph distinguishability is a representational property of the induced feature space (as reflected by Gram-matrix characteristics such as rank, cf. Fig. 4), but it is not sufficient to predict downstream accuracy. Predictive performance also depends on how well kernel-induced similarities align with the labels under the chosen learning procedure, including model selection, regularization, and finite-sample effects, as well as (for quantum features) discretization and finite-shot noise. Hence, increased distinguishability may help, but does not by itself guarantee improved generalization. In particular, two detrimental extremes can arise: if a kernel is overly discriminative on a dataset with no isomorphic pairs, the Gram matrix approaches the identity and provides little usable similarity structure. Conversely, if a kernel assigns nearly identical similarity to all graph pairs, the Gram matrix approaches the all-ones matrix, which is likewise uninformative.